OpenStudy (ckenn):

A spherical fish bowl is half-filled with water. The center of the bowl is C, and the length of segment AB is 12 inches, as shown below. Use Twenty two over seven for pi. Which of the following can be used to calculate the volume of water inside the fish bowl? a) 1/2 (4/3) (22/7) (12^2) (6) b) 1/2 (4/3) (22/7) (12^3) c) 1/2 (4/3) (22/7) (6^2) (12) d) 1/2 (4/3) (22/7) (6^3)

OpenStudy (ckenn):

@Michele_Laino @surjithayer could you help me?

OpenStudy (ckenn):

@TrojanPoem

OpenStudy (trojanpoem):

"as shown below " is there a figure ?

OpenStudy (trojanpoem):

Volume of the water is half the volume of the sphere. Volume of sphere: \[V_{ol}= \frac{ 4 }{ 3 } \pi r^3\] \[r = 0.5 AB = 0.5 * 12 = 6 inch\] \[V_{water} = 0.5 \frac{ 4 }{ 3} \ \frac{22}{7} * (6)^3 = \]

OpenStudy (ckenn):

thank you!!